On July 1, 2019, Cullumber Company purchased new equipment for $85,000. Its estimated useful life was 5 years with a $12,000 salvage value. On December 31, 2022, the company estimated that the equipment’s remaining useful life was 10 years, with a revised salvage value of $5,000.

Compute the revised annual depreciation on December 31, 2022. Revised annual depreciation $__________________

Show your work
Answer #1

Annual depreciation = (85000-12000)/5 = 14600

Net book value on December 31, 2021 = 85000 - 14600*2.5 = $48500

Revised annual depreciation on December 31, 2022 = (48500-5000) / 7.5 YEARS = $5800